OpenStudy (lillian_a):

The position of a particle at any time t, t is (greater than or equal to) 0 is given by s(t)=ln(t+1)+1/8t^2-4t. In which direction is the particle moving when the acceleration is zero?

OpenStudy (anonymous):

Did you mean to write s(t)=...? Construct the equations for velocity and acceleration. v(t)=s'(t) a(t)=v'(t) Then set a(t)=0 and solve for t. Plug this value of t into v(t) to see which way it's going :)

OpenStudy (lillian_a):

\[ s(t)=\ln (t+1)+1/8t^2-4t\]

OpenStudy (anonymous):

Can you tell me what v(t) and a(t) are now?

OpenStudy (anonymous):

I've got to go now so I'm just posting the answer so you can check yours when you're done. \[s(t)=ln(t+1)+\dfrac{1}{8t^2}-4t\]\[v(t)=s'(t)=\dfrac{1}{t+1}-\dfrac{8}{t^2}-4\]\[a(t)=v'(t)=-\dfrac{1}{(t+1)^2} +\dfrac{16}{t^3} \] \[0=-\dfrac{1}{(t+1)^2} +\dfrac{16}{t^3}=-t^3+16(t+1)^2=-t^3+16t^2+32t+16\]\[t\approx17.844\] \[v(17.844)=-3.972057681\]So the particle is moving backwards. I think this is right, but it's been a while since I've done any mechanics.